# On the contraction properties of Sinkhorn semigroups

### v2

@@ 3 Contraction of Markov operators
3.1 Metrics and divergences
3.1.1 Hilbert projective metric
The Hilbert (projective) metric is defined, for any comparable measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$, by the formula
$H(\nu_{1},\nu_{2}):=\log{\left(\left\|\frac{d\nu_{1}}{d\nu_{2}}\right\|\left\|\frac{d\nu_{2}}{d\nu_{1}}\right\|\right)}\geq(\log{3})~\|\nu_{1}-\nu_{2}\|_{\sf tv},$ (31)
where $\|\cdot\|$ and $\|\cdot\|_{\sf tv}$ denote the supremum and total variation norms, respectively, and $H(\nu_{1},\nu_{2})=\infty$ when the measures $\nu_{1},\nu_{2}$ are not comparable. The proof of the r.h.s. estimate in (31) is provided in [atar] (cf. Lemma 1), see also [legland]. For a more thorough discussion on Hilbert projective metrics we refer the reader to the recent review and self contained article on Hilbert projective metrics [cohen].
3.1.2 Integral operator norms
Given some functions $\varphi\in{\cal B}(\mathbb{X})$ such that $\varphi\geq 1/2$, let ${\cal M}_{\varphi}(\mathbb{X})\subset{\cal M}_{b}(\mathbb{X})$ be the subspace of measures $\nu\in{\cal M}_{b}(\mathbb{X})$ equipped with the operator $\varphi$-norm
@@ 3.1.2 Integral operator norms
In the above display $\nu=\nu_{+}-\nu_{-}$ stands for a Hahn-Jordan decomposition of the measure, while $\mbox{\rm osc}_{\varphi}(f)$ stands for the $\varphi$-oscillation of a function $f\in{\cal B}(\mathbb{X})$, which is given by
@@
For a detailed proof of the equivalent formulations in the latter definition (32), we refer to Proposition 8.2.16 in [penev]. The weighted Wasserstein formulation (7) of the $\varphi$-norm is a consequence of Kellerer’s extension of the Kantorovich-Rubinstein theorem [edwards2011kantorovich, kellerer1984duality, kellerer1982duality], see also the recent article [delgerber2025].
@@
Recall that, given an arbitrary Polish space $S$, for any $\mu_{1},\mu_{2}\in{\cal M}_{1}(S)$ and $\epsilon\in]0,1]$ we have
$\|\mu_{1}-\mu_{2}\|_{\sf tv}\leq 1-\epsilon\quad\text{if, and only if,}\quad\exists\nu\in{\cal M}_{1}(S)~:~\mu_{1}\geq\epsilon\nu\quad\mbox{\rm and}\quad\mu_{2}\geq\epsilon\nu.$ (33)
Choosing $\varphi\geq 1$ and $\kappa>0$, for any $\nu\in{\cal M}_{b}(\mathbb{X})$ we also have the norm equivalence formulae
@@
$\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho(x_{1},x_{2}):=\frac{1}{2^{(p-1)_{+}}}~\|b(x_{1})-b(x_{2})\|^{p}$ (36)
as well as the Wasserstein estimate (10).
3.2 Some contraction coefficients
Let $\delta_{x}$ denote the Dirac delta measure located at $x$. Whenever $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$, we denote by $\imath_{x_{1},x_{2}}({\cal K})(y)$ the Radon-Nikodym function
$\imath_{x_{1},x_{2}}({\cal K})(y):=\frac{d\delta_{x_{1}}{\cal K}}{d\delta_{x_{2}}{\cal K}}(y).$ (37)
@@ 3.2 Some contraction coefficients
$\jmath({\cal K})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~\inf_{y\in\mathbb{Y}}~\imath_{x_{1},x_{2}}({\cal K})(y)\quad\text{and}$
$\hbar({\cal K})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\cal K})(y_{1})~\imath_{x_{2},x_{1}}({\cal K})(y_{2})\right)\geq\jmath({\cal K})^{2},$ (38)
@@
$\delta_{x_{1}}{\cal K}(dy_{1})~\delta_{x_{2}}{\cal K}(dy_{2})~\geq\hbar({\cal K})~\delta_{x_{2}}{\cal K}(dy_{1})~\delta_{x_{1}}{\cal K}(dy_{2}).$
@@
$\delta_{x_{1}}{\cal K}(dy_{1})~\geq\hbar({\cal K})~\delta_{x_{2}}{\cal K}(dy_{1}),\quad\text{which readily yields}\quad\jmath({\cal K})\geq\hbar({\cal K})\geq\jmath({\cal K})^{2}.$ (39)
3.2.1 Birkhoff contraction coefficient
The Birkhoff contraction coefficient of a Markov transition $\cal K$} from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
$\raisebox{2.0pt}{$\chi$}_{H}({\cal K}):=\sup\frac{H(\nu_{1}{\cal K},\nu_{2}{\cal K})}{H(\nu_{1},\nu_{2})}=\frac{1-\sqrt{\hbar({\cal K})}}{1+\sqrt{\hbar({\cal K})}},$ (40)
where the supremum is taken over all comparable probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(\mathbb{X})$ such that $H(\nu_{1},\nu_{2})>0$ and the second equality is explicitly derived in Appendix B. Combining (39) and (40) we readily obtain the equivalent inequalities
${\raisebox{2.0pt}{$\chi$}_{H}({\cal K})<1}\Longleftrightarrow\hbar({\cal K})>0\Longleftrightarrow\jmath({\cal K})>0.$
We also emphasize that even on totally bounded state spaces such as $\mathbb{X}=\mathbb{Y}$ there is an interesting twist because the uniform minorization condition $\jmath({\cal K})>0$ is generally not satisfied. We illustrate this assertion with the semi-circle transition on the unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ defined by
${\cal K}(x,dy)=\frac{1}{q(x)}~\sqrt{1-(x-y)^{2}}~1_{]0,1[}(y)~dy$ (41)
for some normalizing constant $1/3\leq q(x)\leq 1$. In this case, choosing $x_{1}=1/(2n)=1-y=1-x_{2}$ in (37) we check that
$0\leq\jmath({\cal K})\leq 3~\sqrt{1-(1-1/n)^{2}}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0\Longrightarrow\jmath({\cal K})=0.$
The entropic transport model (4) on the unit interval ]0,1[]0,1[ with reference transition (41) is discussed in Section 5.4. Note that in that case, the cost function is unbounded.
3.2.2 Dobrushin contraction coefficients
Consider functions $\varphi\in{\cal B}(\mathbb{X})$ and $\psi\in{\cal B}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ for any $(x,y)\in(\mathbb{X}\times\mathbb{Y})$. The $(\varphi,\psi)$-Dobrushin coefficient $\raisebox{2.0pt}{$\chi$}_{\varphi,\psi}({\cal K})$ of a Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by the norm operator
@@ 3.2.2 Dobrushin contraction coefficients
When $(\varphi,\mathbb{X})=(\psi,\mathbb{Y})$ we may also write $\chi_{\varphi}({\cal K})$ instead of $\chi_{\varphi,\varphi}({\cal K})$. Given a Markov transition ${\cal K}_{0}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a Markov transition ${\cal K}_{1}(x,dy)$ from $\mathbb{Y}$ into $\mathbb{X}$ we have
$\chi_{\varphi}({\cal K}_{0}{\cal K}_{1})=\chi_{\varphi,\varphi}({\cal K}_{0}{\cal K}_{1})\leq\chi_{\varphi,\psi}({\cal K}_{0})~\chi_{\psi,\varphi}({\cal K}_{1}).$ (42)
The terminology $\varphi$-Dobrushin coefficient comes from the fact that we recover the standard Dobrushin contraction coefficient $\raisebox{2.0pt}{$\chi$}({\cal K})$ by choosing constant functions. Indeed, if $\varphi=1=\psi$, then
$\raisebox{2.0pt}{$\chi$}_{1,1}({\cal K})=\sup_{\nu_{1},\nu_{2}\,\in{\cal M}_{1}(\mathbb{X})}\frac{\|\nu_{1}{\cal K}-\nu_{2}{\cal K}\|_{\sf tv}}{\|\nu_{1}-\nu_{2}\|_{\sf tv}}=:\raisebox{2.0pt}{$\chi$}({\cal K})=1-\varepsilon({\cal K}),$ (43)
@@
$\varepsilon({\cal K}):=1-\sup_{(x_{1},x_{2})\,\in\mathbb{X}^{2}}{\|(\delta_{x_{1}}-\delta_{x_{2}}){\cal K}\|_{\sf tv}}.$ (44)
@@
Whenever $\delta_{x_{1}}{\cal K}\ll\delta_{x_{2}}{\cal K}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$ using the estimate (39) as well as (33) we readily find that
$\varepsilon({\cal K})\geq\jmath({\cal K})\geq\hbar({\cal K})\quad\mbox{\rm and}\quad\raisebox{2.0pt}{$\chi$}({\cal K})\leq 1-\jmath({\cal K})\leq 1-\hbar({\cal K}).$ (45)
3.2.3 Entropy contraction coefficient
@@
Given a convex function $\Phi$ satisfying (12), the Dobrushin $\Phi$-contraction coefficient $\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K})$ of some Markov transition ${\cal K}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
$\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K}):=\sup_{\nu_{1},\nu_{2}\,\in{\cal M}_{1}(\mathbb{X})}\frac{D_{\Phi}(\nu_{1}{\cal K},\nu_{2}{\cal K})}{D_{\Phi}(\nu_{1},\nu_{2})},$ (46)
@@
where the supremum is taken over all probability measures $\nu_{1},\nu_{2}\in{\cal M}_{1}(S)$ such that $D_{\Phi}(\nu_{1},\nu_{2})>0$. Here, again, the terminology $\Phi$-Dobrushin coefficient comes from the fact that choosing $\Phi(u,v)=\Phi_{0}(u,v)=\frac{1}{2}|u-v|$ we recover the standard Dobrushin contraction coefficient, i.e., $\raisebox{2.0pt}{$\chi$}_{\Phi_{0}}({\cal K})=\raisebox{2.0pt}{$\chi$}({\cal K})$.
For any convex function $\Phi$ satisfying (12) we recall that
$\raisebox{2.0pt}{$\chi$}_{\Phi}({\cal K})\leq\raisebox{2.0pt}{$\chi$}({\cal K}).$ (47)
@@
The estimate (47) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [cohen-0] for finite state spaces and [dm-03] for general measurable spaces). For further details on the Dobrushin contraction coefficient we refer the reader to the article [dob-56], see also [dg-01, dm-03] and Chapter 4 in [dm-04].
4 Sinkhorn contraction properties
4.1 Strong mixing models

### v3

@@ 3. Contraction of Markov operators
3.1. Metrics and divergences
3.1.1. Hilbert projective metric
The Hilbert (projective) metric is defined, for any comparable measures $\nu_{1},\nu_{2}\in{\mathcal{M}}_{1}(\mathbb{X})$, by the formula
(3.1) $H(\nu_{1},\nu_{2}):=\log{\left(\left\|\frac{d\nu_{1}}{d\nu_{2}}\right\|\left\|\frac{d\nu_{2}}{d\nu_{1}}\right\|\right)}\geq(\log{3})~\|\nu_{1}-\nu_{2}\|_{\sf tv},$
where $\|\cdot\|$ and $\|\cdot\|_{\sf tv}$ denote the supremum and total variation norms, respectively, and $H(\nu_{1},\nu_{2})=\infty$ when the measures $\nu_{1},\nu_{2}$ are not comparable. The proof of the r.h.s. estimate in (3.1) is provided in [atar] (cf. Lemma 1), see also [legland]. For a more thorough discussion on Hilbert projective metrics we refer the reader to the recent review and self contained article on Hilbert projective metrics [cohen].
3.1.2. Integral operator norms
Given some functions $\varphi\in{\mathcal{B}}(\mathbb{X})$ such that $\varphi\geq 1/2$, let ${\mathcal{M}}_{\varphi}(\mathbb{X})\subset{\mathcal{M}}_{b}(\mathbb{X})$ be the subspace of measures $\nu\in{\mathcal{M}}_{b}(\mathbb{X})$ equipped with the operator $\varphi$-norm
@@ 3.1.2. Integral operator norms
In the above display $\nu=\nu_{+}-\nu_{-}$ stands for a Hahn-Jordan decomposition of the measure, while $\mbox{\rm osc}_{\varphi}(f)$ stands for the $\varphi$-oscillation of a function $f\in{\mathcal{B}}(\mathbb{X})$, which is given by
@@
For a detailed proof of the equivalent formulations in the latter definition (3.2), we refer to Proposition 8.2.16 in [penev]. The weighted Wasserstein formulation (1.7) of the $\varphi$-norm is a consequence of Kellerer’s extension of the Kantorovich-Rubinstein theorem [edwards2011kantorovich, kellerer1984duality, kellerer1982duality], see also the recent article [delgerber2025].
@@
(3.3) $\|\mu_{1}-\mu_{2}\|_{\sf tv}\leq 1-\epsilon\quad\text{if, and only if,}\quad\exists\nu\in{\mathcal{M}}_{1}(S)~:~\mu_{1}\geq\epsilon\nu\quad\mbox{\rm and}\quad\mu_{2}\geq\epsilon\nu.$
Choosing $\varphi\geq 1$ and $\kappa>0$, for any $\nu\in{\mathcal{M}}_{b}(\mathbb{X})$ we also have the norm equivalence formulae
@@
(3.6) $\varphi(x_{1})+\varphi(x_{2})\geq\varrho(x_{1},x_{2})\quad\mbox{\rm with}\quad\varrho(x_{1},x_{2}):=\frac{1}{2^{(p-1)_{+}}}~\|b(x_{1})-b(x_{2})\|^{p}$
as well as the Wasserstein estimate (1.10).
3.2. Some contraction coefficients
Let $\delta_{x}$ denote the Dirac delta measure located at $x$. Whenever $\delta_{x_{1}}{\mathcal{K}}\ll\delta_{x_{2}}{\mathcal{K}}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$, we denote by $\imath_{x_{1},x_{2}}({\mathcal{K}})(y)$ the Radon-Nikodym function
(3.7) $\imath_{x_{1},x_{2}}({\mathcal{K}})(y):=\frac{d\delta_{x_{1}}{\mathcal{K}}}{d\delta_{x_{2}}{\mathcal{K}}}(y).$
@@ 3.2. Some contraction coefficients
$\jmath({\mathcal{K}})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}~\inf_{y\in\mathbb{Y}}~\imath_{x_{1},x_{2}}({\mathcal{K}})(y)\quad\text{and}$
(3.8) $\hbar({\mathcal{K}})$ $:=$ $\inf_{(x_{1},x_{2})\in\mathbb{X}^{2}}\inf_{(y_{1},y_{2})\in\mathbb{Y}^{2}}\left(\imath_{x_{1},x_{2}}({\mathcal{K}})(y_{1})~\imath_{x_{2},x_{1}}({\mathcal{K}})(y_{2})\right)\geq\jmath({\mathcal{K}})^{2},$
@@
$\delta_{x_{1}}{\mathcal{K}}(dy_{1})~\delta_{x_{2}}{\mathcal{K}}(dy_{2})~\geq\hbar({\mathcal{K}})~\delta_{x_{2}}{\mathcal{K}}(dy_{1})~\delta_{x_{1}}{\mathcal{K}}(dy_{2}).$
@@
(3.9) $\delta_{x_{1}}{\mathcal{K}}(dy_{1})~\geq\hbar({\mathcal{K}})~\delta_{x_{2}}{\mathcal{K}}(dy_{1}),\quad\text{which readily yields}\quad\jmath({\mathcal{K}})\geq\hbar({\mathcal{K}})\geq\jmath({\mathcal{K}})^{2}.$
3.2.1. Birkhoff contraction coefficient
The Birkhoff contraction coefficient of a Markov transition $\mathcal{K}$} from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
(3.10) $\raisebox{2.0pt}{$\chi$}_{H}({\mathcal{K}}):=\sup\frac{H(\nu_{1}{\mathcal{K}},\nu_{2}{\mathcal{K}})}{H(\nu_{1},\nu_{2})}=\frac{1-\sqrt{\hbar({\mathcal{K}})}}{1+\sqrt{\hbar({\mathcal{K}})}},$
where the supremum is taken over all comparable probability measures $\nu_{1},\nu_{2}\in{\mathcal{M}}_{1}(\mathbb{X})$ such that $H(\nu_{1},\nu_{2})>0$ and the second equality is explicitly derived in Appendix B. Combining (3.9) and (3.10) we readily obtain the equivalent inequalities
${\raisebox{2.0pt}{$\chi$}_{H}({\mathcal{K}})<1}\Longleftrightarrow\hbar({\mathcal{K}})>0\Longleftrightarrow\jmath({\mathcal{K}})>0.$
We also emphasize that even on totally bounded state spaces there is an interesting twist because the uniform minorization condition $\jmath({\mathcal{K}})>0$ is generally not satisfied. We illustrate this assertion with the semi-circle transition on the unit interval $\mathbb{X}=]0,1[=\mathbb{Y}$ defined by
(3.11) ${\mathcal{K}}(x,dy)=\frac{1}{q(x)}~\sqrt{1-(x-y)^{2}}~1_{]0,1[}(y)~dy$
for some normalizing constant $1/3\leq q(x)\leq 1$. In this case, choosing $x_{1}=1/(2n)=1-y=1-x_{2}$ in (3.7) we check that
$0\leq\jmath({\mathcal{K}})\leq 3~\sqrt{1-(1-1/n)^{2}}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0\Longrightarrow\jmath({\mathcal{K}})=0.$
The entropic transport model (1.4) on the unit interval ]0,1[]0,1[ with reference transition (3.11) is discussed in Section 5.4. Note that in that case, the cost function is unbounded.
3.2.2. Dobrushin contraction coefficients
Consider functions $\varphi\in{\mathcal{B}}(\mathbb{X})$ and $\psi\in{\mathcal{B}}(\mathbb{Y})$ such that $\varphi(x)\wedge\psi(y)\geq 1$ for any $(x,y)\in(\mathbb{X}\times\mathbb{Y})$. The $(\varphi,\psi)$-Dobrushin coefficient $\raisebox{2.0pt}{$\chi$}_{\varphi,\psi}({\mathcal{K}})$ of a Markov transition ${\mathcal{K}}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by the norm operator
@@ 3.2.2. Dobrushin contraction coefficients
When $(\varphi,\mathbb{X})=(\psi,\mathbb{Y})$ we may also write $\chi_{\varphi}({\mathcal{K}})$ instead of $\chi_{\varphi,\varphi}({\mathcal{K}})$. Given a Markov transition ${\mathcal{K}}_{0}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ and a Markov transition ${\mathcal{K}}_{1}(x,dy)$ from $\mathbb{Y}$ into $\mathbb{X}$ we have
(3.12) $\chi_{\varphi}({\mathcal{K}}_{0}{\mathcal{K}}_{1})=\chi_{\varphi,\varphi}({\mathcal{K}}_{0}{\mathcal{K}}_{1})\leq\chi_{\varphi,\psi}({\mathcal{K}}_{0})~\chi_{\psi,\varphi}({\mathcal{K}}_{1}).$
The terminology $\varphi$-Dobrushin coefficient comes from the fact that we recover the standard Dobrushin contraction coefficient $\raisebox{2.0pt}{$\chi$}({\mathcal{K}})$ by choosing constant functions. Indeed, if $\varphi=1=\psi$, then
(3.13) $\raisebox{2.0pt}{$\chi$}_{1,1}({\mathcal{K}})=\sup_{\nu_{1},\nu_{2}\,\in{\mathcal{M}}_{1}(\mathbb{X})}\frac{\|\nu_{1}{\mathcal{K}}-\nu_{2}{\mathcal{K}}\|_{\sf tv}}{\|\nu_{1}-\nu_{2}\|_{\sf tv}}=:\raisebox{2.0pt}{$\chi$}({\mathcal{K}})=1-\varepsilon({\mathcal{K}}),$
@@
(3.14) $\varepsilon({\mathcal{K}}):=1-\sup_{(x_{1},x_{2})\,\in\mathbb{X}^{2}}{\|(\delta_{x_{1}}-\delta_{x_{2}}){\mathcal{K}}\|_{\sf tv}}.$
@@
Whenever $\delta_{x_{1}}{\mathcal{K}}\ll\delta_{x_{2}}{\mathcal{K}}$ for any $(x_{1},x_{2})\in\mathbb{X}^{2}$ using the estimate (3.9) as well as (3.3) we readily find that
(3.15) $\varepsilon({\mathcal{K}})\geq\jmath({\mathcal{K}})\geq\hbar({\mathcal{K}})\quad\mbox{\rm and}\quad\raisebox{2.0pt}{$\chi$}({\mathcal{K}})\leq 1-\jmath({\mathcal{K}})\leq 1-\hbar({\mathcal{K}}).$
3.2.3. Entropy contraction coefficient
@@
Given a convex function $\Phi$ satisfying (1.12), the Dobrushin $\Phi$-contraction coefficient $\raisebox{2.0pt}{$\chi$}_{\Phi}({\mathcal{K}})$ of some Markov transition ${\mathcal{K}}(x,dy)$ from $\mathbb{X}$ into $\mathbb{Y}$ is defined by
(3.16) $\raisebox{2.0pt}{$\chi$}_{\Phi}({\mathcal{K}}):=\sup_{\nu_{1},\nu_{2}\,\in{\mathcal{M}}_{1}(\mathbb{X})}\frac{D_{\Phi}(\nu_{1}{\mathcal{K}},\nu_{2}{\mathcal{K}})}{D_{\Phi}(\nu_{1},\nu_{2})},$
@@
where the supremum is taken over all probability measures $\nu_{1},\nu_{2}\in{\mathcal{M}}_{1}(S)$ such that $D_{\Phi}(\nu_{1},\nu_{2})>0$. Here, again, the terminology $\Phi$-Dobrushin coefficient comes from the fact that choosing $\Phi(u,v)=\Phi_{0}(u,v)=\frac{1}{2}|u-v|$ we recover the standard Dobrushin contraction coefficient, i.e., $\raisebox{2.0pt}{$\chi$}_{\Phi_{0}}({\mathcal{K}})=\raisebox{2.0pt}{$\chi$}({\mathcal{K}})$. For any convex function $\Phi$ satisfying (1.12) we recall that
(3.17) $\raisebox{2.0pt}{$\chi$}_{\Phi}({\mathcal{K}})\leq\raisebox{2.0pt}{$\chi$}({\mathcal{K}}).$
@@
The estimate (3.17) ensures that the Dobrushin coefficient is a universal contraction coefficient (see [cohen-0] for finite state spaces and [dm-03] for general measurable spaces). For further details on the Dobrushin contraction coefficient we refer the reader to the article [dob-56], see also [dg-01, dm-03] and Chapter 4 in [dm-04].
4. Sinkhorn contraction properties
This section investigates the contraction and stability properties of Sinkhorn iterations and their associated Gibbs-loop semigroups. We first recall the strong-mixing regime, in which uniform positivity of the reference kernel yields Hilbert-metric contractions and exponential convergence. We then move beyond this setting by introducing Lyapunov drift and local minorization conditions tailored to the forward-backward Sinkhorn transitions. These conditions allow us to establish weighted-norm contraction inequalities, from which we derive exponential convergence rates in total variation, weighted norms, Wasserstein distances, and relative entropy. The results provide a unified framework for analyzing Sinkhorn dynamics on non-compact state spaces and for treating models that fall outside the classical strong-mixing assumptions.
4.1. Strong mixing models